what is the interest rate to the nearest tenth of a percent of 20000=15000(1+r)^5?
Answer by LinnW(1048)

divide each side by 15000
20000/15000 = (1+r)^5
4/3 = (1+r)^5
raise each side to the 1/5 power
(4/3)^(1/5) = ((1+r)^5)(1/5)
(4/3)^(1/5) = 1+r
1.0592238 = 1+r
add -1 to each side
0.0592238 = r
5.92238% = r
rounded to the nearest 1/10th
5.9%